OpenStudy (anonymous):

If sin θ=−12/13 with 3π/2<θ<2π , find each of the following: (a) sin(θ−5π/4) (b) cos(θ−11π/6) c) tan(θ+5π/4) If sin θ=−12/13 with 3π/2<θ<2π , find each of the following: (a) sin(θ−5π/4) (b) cos(θ−11π/6) c) tan(θ+5π/4) @Mathematics

OpenStudy (anonymous):

lord we can do this but what a pain. first of all we need \[\cos(\theta)\]

OpenStudy (anonymous):

since you are in quadrant III it is positive, and by pythagoras is it \[\cos(\theta)=\frac{5}{13}\]

OpenStudy (anonymous):

it shuld be negative in quadrant III

OpenStudy (anonymous):

then for \[\sin(\theta-\frac{5\pi}{4})\] use the "subtraction angle formula" \[\sin(\alpha -\beta)=\sin(\alpha)\cos(\beta) -\sin(\beta)\cos(\alpha)\] with \[\sin(\alpha)=-\frac{12}{13}, \cos(\alpha)=\frac{5}{13}, \sin(\beta)=\sin(\frac{5\pi}{4})=-\frac{\sqrt{2}}{2},\cos(\frac{5\pi}{4})=-\frac{\sqrt{2}}{2}\]\]

OpenStudy (anonymous):

my mistake you are in quadrant IV

OpenStudy (anonymous):

plug in the numbers get out the answer

OpenStudy (anonymous):

hw is it quadrant iv?